The sum of the ages of Anup and his father is 100. When Anup is as old as his father now,he will be five times as old as his son Anuj is now.Anuj will be eight years older than Anup is now,when Anup is as old as his father.What are their ages now.Please solve the problem using one variable only,as it is a problem from chapter Linear Equation In one variable

Grade:7

2 Answers

karthika
7 years ago
LetLet age of Anup is x yrs

thus, Father's age= (100-x)

Anuj's age = (100-x)/5

Eq. 1 (Translating 2nd last line#)

(100-x)-x+(100-x)/5=x+8

600-11x=5x+40

16x=540

x =35 yrs

age of Anup is 35 yrs

age of Anuj is 13yrs

age of Anup's father is 65yrs
Arpit
6 years ago
Let anup`s present age be x years His father present age=💯-x year`sAnup present age =💯-x/5 yearsAnup become as old as his father is now after (💯-2x) yearsAfter 💯-2x years we`veAnup`s age =(💯-x/5+💯-2x) years=600-11x/5 yearsAfter (💯-2x)anuj is 8 years older than his father anup is now 600-11x/5=x+816x=560X=35Father=💯-x=💯-35=65Anup`s age =(💯-35/5) =13